OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y= x^(1/2), (1,1)

OpenStudy (freckles):

So you need to differentiate y=x^(1/2) have you done that yet?

OpenStudy (freckles):

\[\text{ the tangent line to } f \text{ at } (a,f(a)) \text{ is } y-f(a)=f'(a)(x-a)\]

OpenStudy (anonymous):

1/2(x^(1/2)) is the derivitive

OpenStudy (anonymous):

Got it, Thank you!

OpenStudy (freckles):

oh you got it ? you need no more help?

OpenStudy (freckles):

that is \[y'=\frac{1}{2 x^{\frac{1}{2}}}\]?

OpenStudy (anonymous):

Thats the same thing I got, thank you